### v1

@@ IV.3 Emulation in ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$, $E_{rel}$, $\lambda$, $\Lambda_{2}$)
The CERPE is examined here by studying both $s$ and $p$ waves. The parameter space is now enlarged to five-dimensional: ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$, $E_{rel}$, $\lambda$, $\Lambda_{2}$). $\Lambda_{2}$ is a non-affine parameter, but we are not concerned with the emulation speed but the accuracy. Extra steps are needed for speed, which will be addressed elsewhere.
@@
As will be illustrated in the results of this subsection, the behaviors of the emulators, when the $H$ and sources are varied, are similar to the behaviors we have seen in the CEE case, except the clear indication of a higher-dimensional solution manifold mapped out by varying more parameters. The $e^{-c\sqrt{N_{\text{eff}}}}$ scaling of the emulation errors is not as clearly seen as in the CEE case, but the general trends are similar.
@@
Let’s start with the $s$-wave case. The training points are sampled using LHS in a 4-dimensional space: $\lambda\in[-3,-1.5]$, $\Lambda_{2}\in[150,250]$ MeV, $E_{rel}\in[0,20]$ MeV, $\mathrm{Re}\,E^{tr}_{\alpha}\in[-20,20]$ MeV, with $\mathrm{Im}\,E^{tr}_{\alpha}$ fixed to either 3 or 30 MeV, as done previously. To test the emulators, $\lambda$, $\Lambda_{2}$, and $E_{rel}$ are sampled randomly 50 times in the same ranges as in sampling the training points, but the ranges for $E$ are more extensive. The $(\mathrm{Re}\,E,\mathrm{Im}\,E)$ values are on a $30\times 30$ evenly-spaced grid spanning from $-50$ to $50$ MeV along both axis to study interpolation and extrapolation in the complex $E$ plane.
@@
Figure 9: For the emulations in the $s$-wave channel in ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$, $E_{rel}$, $\lambda$, $\Lambda_{2}$) space. The plots show the mean of the relative emulation errors of the emulations in the complex $E$ plane. See the text for more details. Note the range of the complex plane is smaller than those in, e.g., Fig. 3.
@@
Fig. 9 shows, at each $E$, the mean of the relative emulation errors, averaged over the sample with the same $E$ but different $(E_{rel},\lambda,\Lambda_{2})$ values. The error pattern is similar to the pattern, e.g., in the left sub-figure of Fig. 3, with a clear sign of moved BCT.
@@
Figure 10: For the same emulators as shown in Fig. 9. But, what is plotted are the distribution of the poles (or emulated spectra) in $\begin{bmatrix}\mathcal{A}\end{bmatrix}$s, at a random testing point with $E_{rel},\lambda,\Lambda_{2}=7.62,2.6,178$ (shown in the title).
@@
Fig. 10 shows the information about the emulated spectra (or the poles in the $\begin{bmatrix}\mathcal{A}\end{bmatrix}$) at a randomly chosen testing point (see the title), in the same way as being presented in Fig. 4. Again, we notice similar emulator behaviors in the CEE cases, including the moved BCTs, the tapered exponential clustering of the poles toward the branch points, how the physical state and the DC states respond differently to changing $\mathrm{Im}\,E^{tr}_{\alpha}$ and $N_{b}$, the existence of the spurious poles when $N_{b}$ is too big, and the density increase of the non-spurious poles when lowering $\mathrm{Im}\,E_{\alpha}$ and increasing $N_{b}$.
@@
Figure 11 shows the emulation errors at the same testing point as in Fig. 10. In the same way as the presentation of Fig. 5, the panels show the emulation errors vs. $\sqrt{N_{\mathrm{eff}}}$ for the properties of the physical (bound) state pole and $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ at the interpolation region and a few points on the real axis. The scaling of errors against $\sqrt{N_{\mathrm{eff}}}$ differs from the scaling in the CEE cases, although the general trend is similar. This reflects that we are working with multivariate approximations, different from the univariate cases with the (near)-optimal rational approximation. The other distinction is in the maximum values of the $N_{\mathrm{eff}}$: for the same $\mathrm{Im}\,E^{tr}_{\alpha}$, $\max(\sqrt{N_{\mathrm{eff}}})$ here is larger than the corresponding maximum in Fig. 5. This is easy to understand as the solution subspace mapped by more parameters has higher dimensions than in the CEE cases.
Figure 12: For the emulation of $p$-wave scattering in the ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$, $E_{rel}$, $\lambda$, $\Lambda_{2}$) space. The plots show the mean of relative emulation errors in the complex $E$ plane in the same way as in Fig. 9. See the text for more details.
Figure 13: For the emulations of the $p$-wave scattering as shown in Fig. 12, but tested at two distinct points (see the titles of the sub-figures) in the parameter space. Each sub-figure shows the properties of the poles in $\begin{bmatrix}\mathcal{A}\end{bmatrix}$, as in Fig. 4. In the left sub-figure, the compressed spectrum has a deep bound state; in the right, there is a resonance state.
Figure 14: For the emulations of the $p$-wave scattering as shown in Fig. 12, but tested at two distinct points (see the titles of the sub-figures) in the parameter space. The presentation follows the one in Fig. 11.
The $p$-wave results, presented in the same fashion as the $s$-wave results, are collected in Figs. 12, 13 and 14. The 5-dimensional parameter space is defined as $\lambda\in[-3,-0.5]$, $\Lambda_{2}\in[150,250]$ MeV, $E_{rel}\in[0,20]$ MeV, and both $\mathrm{Re}\,E$ and $\mathrm{Im}\,E$ $\in[-50,50]$ MeV. The training points are sampled using LHS with the same ranges for $E_{rel}$, $\lambda$, and $\Lambda_{2}$, but a narrower range for $\mathrm{Re}\,E\sim[-20,20]$ MeV. $\mathrm{Im}\,E^{tr}_{\alpha}$ is again fixed to $3$ or $30$ MeV. When testing the averaged emulation performance, $\lambda$, $\Lambda_{2}$, and $E_{rel}$ are sampled randomly 200 times in the parameter space for each $E$ in the complex $E$ plane. We provide results at two test points to probe two types of emulations: one with a bound state and the other with a resonance. The parameter values can be found in the title of each sub-figure.
@@ V Three-body demonstrations
In this section, we further study these emulators in a system of three identical bosons with $s$-wave pairwise and $s$-wave three-body interactions. Of course, the physics become more complex. There could be multiple thresholds, multiple bound states, and a near-threshold resonance state in the spectra, meaning the analytical behaviors of $\mathcal{A}$ and $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ in the complex $E$ plane are closer to realistic systems (such as nucleus) than those in the two-body case. When exploring the $\bm{\theta}$ parameter space, the system changes in nontrivial ways. However, the general observations about the properties and performances of the emulators are consistent with those we saw in the two-body cases, lending further evidence about the robustness of the continuum extraction and emulation methods and our general understanding of the methods.
@@ V.1 Three-body: basics
The Hamiltonian operator in the Schrödinger equation is $H_{sr}=H_{0}+V$ [^18] Here, we intend to make a distinction between the Hamiltonian in the Schrödinger equation and the effective Hamiltonian to be used later for studying emulators within the Faddeev framework., with $H_{0}$ as the kinetic energy operator in the center of mass frame, $V\equiv\sum_{i=1}^{4}V_{i}$, $V_{i=1,2,3}$ as the pair-wise interaction with $i$th particle as the spectator, and $V_{4}$ as the three-body force. The notations in Ref. Glöckle (1983) are followed loosely here. All the interactions are separable, with Gaussian-like form factors. The two-body interaction is the same as the interaction discussed in Sec. IV. The system has been studied previously in Ref. Zhang and Furnstahl (2022) for a different purpose.
We work in the Faddeev framework to solve the Schrödinger equation. Therefore, we deal with the Faddeev equations and their solutions, known as Faddeev components, in emulator development and high-fidelity benchmark calculations. It will be interesting to study the equivalence between the current emulation approach and the approach directly based on the Schrödinger equation in the future.
@@ V.1.1 Three-body interaction
Similar to Refs. Phillips (1966); Zhang and Furnstahl (2022), we introduce separable potential for the three-body force with a coupling $\lambda_{4}$: $V_{4}=\lambda_{4}|g_{4}\rangle\langle g_{4}|$. The corresponding $T$-matrix, $t_{4}$, defined with only the three-body interaction, is
@@
The form factor $|g_{4}\rangle$ takes a Gaussian form Zhang and Furnstahl (2022):
@@
As in Ref. Glöckle (1983), $\mu^{1}=2M_{N}/3$ and $\mu_{1}=M_{N}/2$ are the reduced mass between particle and dimer and between particle and particle, respectively; $P_{1}$ and $q_{1}$ are the corresponding relative momenta, with particle-1 as the spectator. The momentum pair with other spectators are $P_{2},q_{2}$ and $P_{3},q_{3}$. Since the kinetic energy and $E_{4}$ are invariant when changing the momenta sets, we have $\langle{P}_{2},{q}_{2}|g_{4}\rangle=\langle{P}_{3},{q}_{3}|g_{4}\rangle=\langle{P}_{1},{q}_{1}|g_{4}\rangle$, which is a property we use to simplify the calculation.
@@ V.1.2 Faddeev equations
Our interest is mainly the elastic scattering between particle and dimer, both below and above the three-body break-up threshold. As a preparation, suppose $|\phi_{1}\rangle\equiv|P_{rel}\rangle|\varphi_{B}\rangle$, which is the product of the bound state of dimer-23 and a plane wave state for the relative motion between particle-1 and the dimer-23. Note $|P_{rel}\rangle$ is a $s$-wave state. The associated many-body total energy of $|\phi_{1}\rangle$ is $E={P_{rel}^{2}}/{2\mu^{1}}-B_{2}$. A variable for the relative energy between a particle and a dimer $E_{rel}={P_{rel}^{2}}/{2\mu^{1}}$ is also used later. For the separable potential, the dimer bound state $|\varphi_{B}\rangle=\xi G_{0,2b}(-B_{2})|g\rangle$ with $G_{0,2b}(-B_{2})$, the free two-body Green’s function evaluated at two-body energy $e_{2}=-B_{2}$ ($B_{2}$ as the dimer binding energy); $\xi$ is a factor properly normalizing the bound state.
@@
and $V^{i}\equiv V-V_{i}$. The Lippmann-Schwinger equation for the channel of scattering beween particle-1 and dimer-23 can be recast as Glöckle (1983)
@@
$|\Psi_{1}^{+}\rangle$ is the full scattering wave function of this channel labeled with subscript $1$. Now, we introduce four different Faddeev components and specifically separate the incoming wave from these components:
$|\psi_{1,1}\rangle$ $=G_{0}V_{1}|\Psi_{1}^{+}\rangle-|\phi_{1}\rangle$
@@
We then arrive at the Faddeev equations for the channel-1 scattering:
@@
We symmetrize the above equations for the identical boson system by summing up the other two channels with particle-2 and -3 as the spectator, respectively. Four Faddeev components in the boson-dimer scattering channel then turn into
@@
Note $|\psi_{2}\rangle$ and $|\psi_{3}\rangle$ can be computed by applying proper permutation operators on $|\psi_{1}\rangle$. For example, $|\psi_{2}\rangle+|\psi_{3}\rangle=\mathbb{P}|\psi_{1}\rangle$, with $\mathbb{P}$ as the sum of two permutation operators Glöckle (1983). Therefore, we focus on the $|\psi_{1}\rangle$ and $|\psi_{4}\rangle$ components. Their Faddeev equations become
@@
$|{S}\rangle\equiv\begin{pmatrix}V_{1}\mathbb{P}|\phi_{1}\rangle\\ 3V_{4}|\phi_{1}\rangle\end{pmatrix}\equiv\begin{pmatrix}|{S}_{1}\rangle\\ |{S}_{4}\rangle\end{pmatrix}\ .$ (55)
Now $H$—not $H_{sr}$—is the effective Hamiltonian operator working in the space of $|\psi_{1}\rangle$ and $|\psi_{4}\rangle$; $|\Psi\rangle$ in these equations is not the whole wave function $|\Psi^{+}\rangle$. To simplify the equations, we take advantage of the fact that $g_{4}(P,q)$ is invariant with respect to particle permutation (see Sec. V.1.1) so that $V_{4}(1+\mathbb{P})=3V_{4}$.
Knowing the full scattering wave function $|\Psi^{+}\rangle$ and its expression in terms of Faddeev components via Eq. (52), we can express the corresponding scattering amplitude as Glöckle (1983)
@@
$\equiv T_{Born}+T_{nonBorn}\ ,$ (56)
@@
$|\tilde{{S}}\rangle\equiv\begin{pmatrix}\bigg{\{}(1+\mathbb{P})V_{1}\mathbb{P}+3V_{4}\bigg{\}}|\phi_{1}\rangle\\ \left(V_{1}\mathbb{P}+V_{4}\right)|\phi_{1}\rangle\end{pmatrix}\equiv\begin{pmatrix}|\tilde{{S}}_{1}\rangle\\ |\tilde{{S}}_{4}\rangle\end{pmatrix}\ .$ (57)
The $T_{Born}$ terms is explicitly presented in Appendix C.1.
Since we are interested in the overlaps between $|\Psi\rangle$ and the source $|\tilde{{S}}\rangle$ for computing $T_{nonBorn}$, we introduce the adjoint equation to develop emulators (see Sec. III.1)

### v2

@@ IV.3 Emulation in ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$, $E_{\mathrm{rel}}$, $\lambda$, $\Lambda_{2}$)
The CERPE is examined here by studying both $s$ and $p$ waves. The parameter space is now enlarged to five-dimensional: ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$, $E_{\mathrm{rel}}$, $\lambda$, $\Lambda_{2}$). $\Lambda_{2}$ is a non-affine parameter, but I am not concerned with the emulation speed but the accuracy. Extra steps are needed for speed, which will be addressed elsewhere.
@@
As will be illustrated in the results of this subsection, the behaviors of the emulators, when the $H$ and sources are varied, are similar to the behaviors seen in the CEE case, except the clear indication of a higher-dimensional solution manifold mapped out by varying more parameters and a different $N_{\mathrm{eff}}$ dependence of the error from $e^{-c\sqrt{N_{\text{eff}}}}$ scaling seen in the CEE case.
@@
I start with the $s$-wave case. The training points are sampled using LHS in a 4-dimensional space: $\lambda\in[-3,-1.5]$, $\Lambda_{2}\in[150,250]$ MeV, $E_{\mathrm{rel}}\in[0,20]$ MeV, $\mathrm{Re}\,E^{\mathrm{tr}}_{\alpha}\in[-20,20]$ MeV, with $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ fixed to either 3 or 30 MeV, as done previously. To test the emulators, $\lambda$, $\Lambda_{2}$, and $E_{\mathrm{rel}}$ are sampled randomly 50 times in the same ranges as in sampling the training points, but the ranges for $E$ are more extensive. The $(\mathrm{Re}\,E,\mathrm{Im}\,E)$ values are on a $30\times 30$ evenly-spaced grid spanning from 50-50 to $50$ MeV along both axis to study interpolation and extrapolation in the complex $E$ plane.
@@
Figure 8: For the emulations in the $s$-wave channel in ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$, $E_{\mathrm{rel}}$, $\lambda$, $\Lambda_{2}$) space. The plots show the mean of the relative emulation errors of the emulations in the complex $E$ plane. See the text for more details. Note the range of the complex plane is smaller than those in, e.g., Fig. 3.
@@
Fig. 8 shows, at each $E$, the mean of the relative emulation errors, averaged over the sample with the same $E$ but different $(E_{\mathrm{rel}},\lambda,\Lambda_{2})$ values. The error pattern is similar to the pattern, e.g., in the left sub-figure of Fig. 3, with a clear sign of moved BCT.
@@
Figure 9: For the same emulators as shown in Fig. 8. But, what is plotted are the distribution of the poles (or emulated spectra) in $\begin{bmatrix}\mathcal{A}\end{bmatrix}$s, at a random testing point with $\lambda,\Lambda_{2},E_{\mathrm{rel}}=-2.60,177.96,7.62$ (shown in the title).
@@
Fig. 9 shows the information about the emulated spectra (or the poles in the $\begin{bmatrix}\mathcal{A}\end{bmatrix}$) at a randomly chosen testing point (see the title), in the same way as being presented in Fig. 4. Again, one notices similar emulator behaviors in the CEE cases, including the moved BCTs, the tapered exponential clustering of the poles toward the branch points, how the physical state and the DC states respond differently to changing $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ and $N_{b}$, the existence of the spurious poles when $N_{b}$ is too big, and the density increase of the non-spurious poles when lowering $\mathrm{Im}\,E_{\alpha}$ and increasing $N_{b}$.
@@
Figure 10 shows the emulation errors at the same testing point as in Fig. 9. In the same way as the presentation of Fig. 5, the panels show the emulation errors vs. $\sqrt{N_{\mathrm{eff}}}$ for the properties of the physical (bound) state pole and $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ at the interpolation region and a few points on the real axis. The scaling of errors against $\sqrt{N_{\mathrm{eff}}}$ differs from the scaling in the CEE cases. This reflects that I am working with multivariate approximations, different from the univariate cases with the (near-)optimal rational approximation. However, do note that the errors still decrease in a certain exponential fashion with $\sqrt{N_{\mathrm{eff}}}$. As the other distinction, the maximum $N_{\mathrm{eff}}$ value here is larger than that in Fig. 5 with the same $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ (see the symbol densities of the insets of the two figures). For example, when $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}=3$ MeV, $N_{\mathrm{eff}}=13$ and $19$ for the CEE and CERPE cases respectively. This is also expected, since the solution subspaces here have higher dimensionalities than those in the CEE case.
Figure 11: For the emulation of $p$-wave scattering in the ($\mathrm{Re}\,E$, $\mathrm{Im}\,E$, $E_{\mathrm{rel}}$, $\lambda$, $\Lambda_{2}$) space. The plots show the mean of relative emulation errors in the complex $E$ plane in the same way as in Fig. 8. See the text for more details.
The $p$-wave results, presented in the same fashion as the $s$-wave results, are collected in Fig. 11 and additional figures in the SM [110]. The 5-dimensional parameter space is defined as $\lambda\in[-3,-0.5]$, $\Lambda_{2}\in[150,250]$ MeV, $E_{\mathrm{rel}}\in[0,20]$ MeV, and both $\mathrm{Re}\,E$ and $\mathrm{Im}\,E$ $\in[-50,50]$ MeV. The training points are sampled using LHS with the same ranges for $E_{\mathrm{rel}}$, $\lambda$, and $\Lambda_{2}$, but a narrower range for $\mathrm{Re}\,E\sim[-20,20]$ MeV. $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ is again fixed to $3$ or $30$ MeV. When testing the averaged emulation performance, $\lambda$, $\Lambda_{2}$, and $E_{\mathrm{rel}}$ are sampled randomly 200 times in the parameter space for each $E$ in the complex $E$ plane. I provide results at two test points in the SM [110] to probe two types of emulations: one with a bound state and the other with a resonance. The parameter values can be found in the title of each sub-figure therein.
@@ V Three-body demonstrations
In this section, I further study these emulators in a system of three identical bosons with $s$-wave pairwise and $s$-wave three-body interactions. Of course, the physics become more complex. There could be multiple thresholds, multiple bound states, and a near-threshold resonance state in the spectra, meaning the analytical behaviors of $\mathcal{A}$ and $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ in the complex $E$ plane are closer to realistic systems (such as nucleus) than those in the two-body case. When exploring the $\bm{\theta}$ parameter space, the system changes in nontrivial ways. However, the general observations about the properties and performances of the emulators are consistent with those seen in the two-body cases, lending further evidence about the robustness of the continuum extraction and emulation methods and the general understanding of the methods.
@@ V.1 Three-body: basics
The Hamiltonian operator in the Schrödinger equation is $H_{sr}=H_{0}+V$, with $H_{0}$ as the kinetic energy operator in the center of mass frame, $V\equiv\sum_{i=1}^{4}V_{i}$, $V_{i=1,2,3}$ as the pair-wise interaction with $i$th particle as the spectator, and $V_{4}$ as the three-body force. The notations in Ref. [49] are followed loosely here. All the interactions are separable, with Gaussian-like form factors. The two-body interaction is the same as the interaction discussed in Sec. IV. The system has been studied previously in Ref. [40] for a different purpose.
I work in the Faddeev framework to solve the Schrödinger equation. Therefore, I deal with the Faddeev equations and their solutions, known as Faddeev components, in emulator development and high-fidelity benchmark calculations. It will be interesting to study the equivalence between the current emulation approach and the approach directly based on the Schrödinger equation in the future.
@@ V.1.1 Three-body interaction
Similar to Refs. [113, 40], I introduce separable potential for the three-body force with a coupling $\lambda_{4}$: $V_{4}=\lambda_{4}|g_{4}\rangle\langle g_{4}|$. The corresponding $T$-matrix, $t_{4}$, defined with only the three-body interaction, is
@@
The form factor $|g_{4}\rangle$ takes a Gaussian form [40]:
@@
As in Ref. [49], $\mu^{1}=2M_{N}/3$ and $\mu_{1}=M_{N}/2$ are the reduced mass between particle and dimer and between particle and particle, respectively; $P_{1}$ and $q_{1}$ are the corresponding relative momenta, with particle-1 as the spectator. The momentum pair with other spectators are $P_{2},q_{2}$ and $P_{3},q_{3}$. Since the kinetic energy and $E_{4}$ are invariant when changing the momenta sets, I have $\langle{P}_{2},{q}_{2}|g_{4}\rangle=\langle{P}_{3},{q}_{3}|g_{4}\rangle=\langle{P}_{1},{q}_{1}|g_{4}\rangle$, which is a property I use to simplify the calculation.
@@ V.1.2 Faddeev equations
My interest is mainly the elastic scattering between particle and dimer, both below and above the dimer break-up threshold. As a preparation, suppose $|\phi_{1}\rangle\equiv|P_{\mathrm{rel}}\rangle|\varphi_{B}\rangle$, which is the product of the bound state of dimer-23 and a plane wave state for the relative motion between particle-1 and the dimer-23. Note $|P_{\mathrm{rel}}\rangle$ is a $s$-wave state. The associated many-body total energy of $|\phi_{1}\rangle$ is $E={P_{\mathrm{rel}}^{2}}/{2\mu^{1}}-B_{2}$. A variable for the relative energy between a particle and a dimer $E_{\mathrm{rel}}={P_{\mathrm{rel}}^{2}}/{2\mu^{1}}$ is also used later. For the separable potential, the dimer bound state $|\varphi_{B}\rangle=\xi G_{0,2b}(-B_{2})|g\rangle$ with $G_{0,2b}(-B_{2})$, the free two-body Green’s function evaluated at two-body energy $e_{2}=-B_{2}$ ($B_{2}$ as the dimer binding energy); $\xi$ is a factor properly normalizing the bound state.
@@
and $V^{i}\equiv V-V_{i}$. The Lippmann-Schwinger equation for the channel of scattering beween particle-1 and dimer-23 can be recast as [49]
@@
$|\Psi_{1}^{+}\rangle$ is the full scattering wave function of this channel labeled with subscript $1$. Now, I introduce four different Faddeev components and specifically separate the incoming wave from these components:
$|\psi_{1,1}\rangle$ $=G_{0}V_{1}|\Psi_{1}^{+}\rangle-|\phi_{1}\rangle\,$,
@@
I then arrive at the Faddeev equations for the channel-1 scattering:
@@
The above equations are then symmetrized for the identical boson system by summing up the other two channels with particle-2 and -3 as the spectator, respectively. Four Faddeev components in the boson-dimer scattering channel then turn into
@@
Note $|\psi_{2}\rangle$ and $|\psi_{3}\rangle$ can be computed by applying proper permutation operators on $|\psi_{1}\rangle$. For example, $|\psi_{2}\rangle+|\psi_{3}\rangle=\mathbb{P}|\psi_{1}\rangle$, with $\mathbb{P}$ as the sum of two permutation operators [49]. Therefore, I focus on the $|\psi_{1}\rangle$ and $|\psi_{4}\rangle$ components. Their Faddeev equations become
@@
$|{S}\rangle\equiv\begin{pmatrix}V_{1}\mathbb{P}|\phi_{1}\rangle\\ 3V_{4}|\phi_{1}\rangle\end{pmatrix}\equiv\begin{pmatrix}|{S}_{1}\rangle\\ |{S}_{4}\rangle\end{pmatrix}\$. (58)
Now HH—not $H_{sr}$—is the effective Hamiltonian operator working in the space of $|\psi_{1}\rangle$ and $|\psi_{4}\rangle$; $|\Psi\rangle$ in these equations is not the whole wave function $|\Psi^{+}\rangle$. To simplify the equations, I take advantage of the fact that $g_{4}(P,q)$ is invariant with respect to particle permutation (see Sec. V.1.1) so that $V_{4}(1+\mathbb{P})=3V_{4}$.
Knowing the full scattering wave function $|\Psi^{+}\rangle$ and its expression in terms of Faddeev components via Eq. (55), the corresponding scattering amplitude can be expressed as [49]
@@
$\equiv T_{\mathrm{Born}}+T_{\mathrm{nonBorn}}\ ,$ (59)
@@
$|\tilde{{S}}\rangle\equiv\begin{pmatrix}\bigg\{(1+\mathbb{P})V_{1}\mathbb{P}+3V_{4}\bigg\}|\phi_{1}\rangle\\ \left(V_{1}\mathbb{P}+V_{4}\right)|\phi_{1}\rangle\end{pmatrix}\equiv\begin{pmatrix}|\tilde{{S}}_{1}\rangle\\ |\tilde{{S}}_{4}\rangle\end{pmatrix}\$. (60)
The $T_{\mathrm{Born}}$ terms is explicitly presented in Appendix C.1.
Since I am interested in the overlaps between $|\Psi\rangle$ and the source $|\tilde{{S}}\rangle$ for computing $T_{\mathrm{nonBorn}}$, I introduce the adjoint equation to develop emulators (see Sec. III.1)